OpenStudy (anonymous):

What is the factoring form of 4x^2 + 12x + 5? Will fan and medal.

jagr2713 (jagr2713):

Well how would you factor this?

OpenStudy (anonymous):

break it down?

jagr2713 (jagr2713):

Yea what number goes into 5?

OpenStudy (anonymous):

2.5?

jagr2713 (jagr2713):

Well does 2.5 added equal to 12?

OpenStudy (anonymous):

no

jagr2713 (jagr2713):

so what do we have to use? it starts with an A

OpenStudy (aishalicious1):

(2x+5) (2x+1) sorry its late

OpenStudy (aishalicious1):

Because it equals 4x^2 +12x+5

OpenStudy (jdcap):

\[4x^2+12x+5 \\ \frac{(4x)^2+12(4x)+20}{4}\\\frac{(4x+10)(4x+2)}{4}\\ \frac{(2x+5)(4x+2)}{2} \\ (2x+5)(2x+1)\]